OpenStudy (anonymous):

Please help mee ♥ A solid metal cylinder with a 4-in. radius and a 10-in. altitude is melted and recast into solid right circular cones each with a 1-in. radius and a 2-in. altitude. The number of cones formed is 106 2/3 120 240

OpenStudy (nurali):

vol of cyl = πr^2h = 160π vol of cone = π/3r^2h = 2π/3 160π / (2π/3) = 240 cones

OpenStudy (anonymous):

Thank you (:

OpenStudy (nurali):

Anytime.